a total of 104 adults and children attend a school concert. the number of children attending was 5 more then twice the number of adults who attended. how many adults attended the concert.
Answer by addingup(3677)

children = adults*2+5 We will use this value for c in the next equation:
c+a = 104 Substitute for children:
2a+5+a = 104 Add a on left and subtract 5 on both sides:
3a = 99
a= 99/3
a= 33 you have 33 adults and 104-33 = 71 children